# WebGen-Bench: Evaluating LLMs on Generating Interactive and Functional Websites from Scratch

### v1

@@ 3.3 Evaluation of Website Appearance
Apart from the fulfillment of the functionality and appearance constraints in the instructions, another important aspect of website generation is the level of relevance, harmony, and aesthetics of the webpage. To conduct a quantitative analysis of this aspect, we designed a set of detailed metrics, ranging from the success of rendering and the relevance of the content to the harmony of the layout and the modernness of the design. We then place the metrics in a prompt, asking GPT-4o to grade the appearance of the website with a score ranging from 1 to 5 (the higher the better), as demonstrated in the middle part of Fig. 1 (b). The prompt is shown in Fig. 13. Examples are presented in Appendix. M.
3.4 Analysis of Dataset Attributes
(a) Instruction-string lengths
(b) Test case numbers
Figure 2: Distributions of instruction lengths (left) and test case numbers per instruction (right).
Table 3: Statistics for the instruction string lengths and the test case numbers in our dataset.
Statistic Minimum Maximum Median Average
Instruction Length (chars) 324 876 483 496.8
Test Case Number 4 11 6 6.4
In this section, we analyze the distribution of instruction lengths and the number of test cases per instruction. The corresponding plots are shown in Fig.2, and the minimum, maximum, median, and average values are summarized in Tab.3.
As depicted in Fig. 2,(a), most website-generation instructions contain between 400 and 600 characters, with a median length of 483 and an average length of 496.8. These relatively long prompts add considerable complexity, posing a meaningful challenge to the agents under evaluation. Fig. 2,(b) indicates that most instructions are associated with five to seven test cases. The median and average numbers of test cases are 6 and 6.4, respectively. Because each test case corresponds to a distinct requirement in the instruction, these statistics confirm that every instruction encompasses a sufficient set of functional and appearance requirements.
@@ 4.1 Experimental Setup
Frameworks. We evaluate three popular code-agent frameworks: Bolt.diy [30], OpenHands [33], and Aider [1]. Bolt.diy is the open-source version of Bolt.new [^6] https://bolt.new, a browser-based framework for generating and previewing web applications. It provides a user interface and a Linux-like WebContainer environment that can execute code. It first prompts the model to decide which frontend and backend frameworks to use (such as Vite, React, Remix, etc.), then imports the basic template and builds upon it. OpenHands is a platform for AI-powered software development agents. For OpenHands, we pair it with CodeActAgent to evaluate it on our benchmark. The adapted instruction is presented in Appendix I. Aider is a terminal-based AI programming framework that natively supports many popular programming languages, including Python, JavaScript, PHP, HTML, CSS, and more. Aider constructs a map of the entire codebase, which helps it function well in larger projects. We use the adapted instruction in Appendix J to generate websites with Aider.
Models. We first evaluate the three frameworks on DeepSeek-V3 [17], a model that is both performant and cost-effective. We then evaluate several strong general-purpose proprietary and open-source LLMs—including Claude-3.5-Sonnet [3], DeepSeek-R1 [10], GPT-4o [14], o3-mini [25], Qwen2.5-Coder-32B [13], and Qwen2.5-72B-Instruct [37]—on the best-performing framework, Bolt.diy. We do not test general-purpose models smaller than Qwen2.5-Coder-32B, as we observe that such models often fail to follow the specified output format and therefore cannot generate valid websites.
Training Details. To validate the effectiveness of our training set, we selectively generated Bolt.diy trajectories for a subset of 2K instructions from WebGen-Instruct using DeepSeek-V3. Using rejection sampling [40], we retained only the trajectories whose corresponding websites achieved an appearance score greater than or equal to 3, resulting in 600 trajectories. This filtering ensures that the remaining generated websites are relevant to the instructions and do not exhibit major rendering issues. We then fine-tuned Qwen2.5-Coder-Instruct models of sizes 7B, 14B, and 32B for 2 epochs, with a learning rate of 4e-5 and a batch size of 32. The 7B, 14B, and 32B models were trained on 8, 16, and 32 A800 GPUs, respectively. This fine-tuning process yields a family of models specialized in website generation, which we name WebGen-LM.
@@ 4.2 Experimental Results
Main Results. Based on the experimental results, we make the following observations: (1) As shown in Tab. 5, WebGen-LM-32B achieves the highest accuracy of 38.2%, surpassing the best proprietary model, DeepSeek-R1, by 10.4%, demonstrating the effectiveness of our training set and the rejection-sampling process. (2) Bolt.diy with DeepSeek-R1 as the engine achieves the highest accuracy among general LLMs at 27.8%, closely followed by Claude-3.5-Sonnet with an accuracy of 26.4%. This indicates that the best-performing models are still far from saturating WebGen-Bench, highlighting that our benchmark remains challenging for current LLMs and agent frameworks. (3) Smaller general open-source models, such as Qwen2.5-Coder-32B and Qwen2.5-72B-Instruct, show a significant performance gap compared to proprietary models. (4) In terms of appearance scores, Bolt.diy with Claude-3.5-Sonnet achieves the best performance of 3.0. The appearance score exhibits a loose correlation with accuracy, as functional webpages typically do not suffer from major rendering issues. To better understand the statistical characteristics of the generated websites, we analyzed the file count and line count in the generated codebases, as detailed in Appendix K.
Categorical Results. Apart from the three main instruction categories (shown in Tab. 2), we also classify the test cases into three primary categories based on what they are intended to evaluate: Functional Testing, Data Display Testing, and Design Validation Testing. Detailed definitions and statistics for these categories are provided in Fig. 18 and Tab. 9 in Appendix L. As shown in Tab. 5, among the different categories of test cases, Design Validation Testing achieves the highest accuracy in most cases, while Functional Testing generally yields lower accuracy. Among instruction categories, Content Presentation consistently demonstrates the highest accuracies. This indicates that superficial aspects, such as color themes, are easier to implement than deeper internal functionalities.
@@ 4.3 Ablation Studies
Analysis of the Accuracy of UI Agent Testing Results. To analyze the accuracy of the UI agent testing process, we manually examined three sets of testing results on Bolt.diy. We select the results of Claude-3.5-Sonnet, DeepSeek-R1, and DeepSeek-V3 as the accuracies of these three models are high and are close to each other. The manual testing results serve as the ground truth and require precision; therefore, three human testers independently annotated the results and we assessed the consistency of their annotations. If the annotations of a test case are inconsistent, a fourth human tester is tasked with re-examining the test case and the inconsistent annotations to decide on a final annotation. We present the results of manual testing in Tab. 6. The Alignment Rate is computed with Alignment $\text{Alignment Rate}=\frac{N_{\text{Manual}=\text{Agent}}}{N_{\text{total}}}\times 100\%$, where $N_{\text{Manual}=\text{Agent}}$ denotes the number of test cases where the agent-generated result aligns with the manually-annotated result.
Figure 4: The distribution of the task errors.
@@
Analysis of Errors in WebGen-Bench Tasks. We analyze the errors and flaws that occur in web generation pipelines using mainstream LLMs such as DeepSeek-V3, as illustrated in Fig. 20. A detailed explanation of the error types can be found in Appendix N. In addition, we present statistics on the distribution of error types for each task in the test set, as shown in Fig. 4. Currently, more than half of the task errors are due to failures in launching a web page or in modifying a template. These results highlight significant potential for future research focused on improving the success rate of web page initialization and template adaptation in generated websites.
@@ Appendix D Details of the Decontamination Process
We then inspect whether the remaining training samples contain instructions that are semantic duplicates of the instructions in the testing. For each testing instruction, we retrieve the top-3 training instructions with the highest cosine similarity, and manually inspect them for semantic duplication. We found that the retrieved training samples are all completely different from the testing samples, proving that the final training set is not contaminated. The first three samples in WebGen-Bench and their top matches in the training set are shown in Fig. 7, Fig. 8, and Fig. 9, respectively. The matches are completely different from the test samples. Fig. 6 shows the distribution of the cosine similarity between the test set and the training set. The cosine similarity is gathered around 0.2 to 0.3, which is relatively low.
Figure 6: The distribution of the cosine similarity between the test set and the training set.
@@ Appendix I Prompt for Adapting OpenHands Paired with CodeActAgent for WebGen-Bench Evaluation
Appendix K Analysis of Average File Count and Average Line Count
Fig. 16 reports the average file and line counts produced by each model, while Fig. 17(a) and Fig. 17(b) show the distributions of file and line counts generated specifically by WebGen-LM-32B. As shown in Fig. 16, GPT-4o, o3-mini, Qwen2.5-72B-Instruct, and Qwen2.5-Coder-32B-Instruct exhibit high average file counts relative to their average line counts, yet their overall performance remains relatively low. One plausible explanation is that, although these models create many files, the files are poorly organized and each contains too little code to support a complete website.
By contrast, the WebGen-LM models generate more lines of code without disproportionately increasing the number of files. Their average line counts all exceed those of DeepSeek-V3, the teacher model used during distillation—an effect that can partly be attributed to the use of rejection sampling. For every WebGen-LM variant, both the file count and the line count rise consistently with model size, indicating that the generated websites become increasingly comprehensive and complex as model scale grows.
As shown in Fig. 17, most samples contain between 4 and 10 files, while their line counts are largely concentrated between 400 and 500. Only two samples include more than 15 files.
Figure 16: The average file and line counts of each model using Bolt.diy as the framework.
(a) File-count distribution.
(b) Line-count distribution.
Figure 17: Distributions of the number of files and lines produced by WebGen-LM-32B.
Appendix L Test Case Categories

### v2

@@ 3.3 Evaluation of Website Appearance
Figure 2: The examples of website screenshots at each appearance score.
Apart from the fulfillment of the functionality and appearance constraints in the instructions, another important aspect of website generation is the level of relevance, harmony, and aesthetics of the webpage. To conduct a quantitative analysis of this aspect, we designed a set of detailed metrics, ranging from the success of rendering and the relevance of the content to the harmony of the layout and the modernness of the design. We then place the metrics in a prompt, asking GPT-4o to grade the appearance of the website with a score ranging from 1 to 5 (the higher the better), as demonstrated in the middle part of Fig. 1 (b). The prompt is shown in Fig. 11. Examples of websites at different appearance scores are shown in Fig. 2, and more examples are presented in Fig. 15 of Appendix. M.
@@ 4.1 Experimental Setup
Frameworks. We evaluate three popular code-agent frameworks: Bolt.diy (stackblitzlabs2024bolt, ), OpenHands (wang2024openhands, ), and Aider (aiderai2024aider, ). Bolt.diy is the open-source version of Bolt.new [^12] https://bolt.new, a browser-based framework for generating and previewing web applications. It provides a user interface and a Linux-like WebContainer environment that can execute code. It first prompts the model to decide which frontend and backend frameworks to use (such as Vite, React, Remix, etc.), then imports the basic template and builds upon it. OpenHands is a platform for AI-powered software development agents. For OpenHands, we pair it with CodeActAgent to evaluate it on our benchmark. The adapted instruction is presented in Appendix I. Aider is a terminal-based AI programming framework that natively supports many popular programming languages, including Python, JavaScript, PHP, HTML, CSS, and more. Aider constructs a map of the entire codebase, which helps it function well in larger projects. We use the adapted instruction in Appendix J to generate websites with Aider.
Models. We first evaluate the three frameworks on DeepSeek-V3 (liu2024deepseek, ), Claude-3.5-Sonnet (anthropic2024claude, ), and DeepSeek-R1 (guo2025deepseek, ). We then evaluate several strong general-purpose proprietary and open-source LLMs—including Claude-3.5-Sonnet (anthropic2024claude, ), DeepSeek-R1 (guo2025deepseek, ), GPT-4o (hurst2024gpt, ), O3-mini (openai2025o3, ), Qwen2.5-Coder-32B (hui2024qwen2, ), and Qwen2.5-72B-Instruct (yang2024qwen2, )—on the best-performing framework, Bolt.diy. We do not test general-purpose models smaller than Qwen2.5-Coder-32B, as we observe that such models often fail to follow the specified output format and therefore cannot generate valid websites.
Training Details. To validate the effectiveness of our training set, we selectively generated Bolt.diy trajectories for a subset of 2K instructions from WebGen-Instruct using DeepSeek-V3. Using rejection sampling (yuan2023scaling, ), we retained only the trajectories whose corresponding websites achieved an appearance score greater than or equal to 3, resulting in 600 trajectories. This filtering ensures that the remaining generated websites are relevant to the instructions and do not exhibit major rendering issues. We then fine-tuned Qwen2.5-Coder-Instruct models of sizes 7B, 14B, and 32B for 2 epochs, with a learning rate of 4e-5 and a batch size of 32. The 7B, 14B, and 32B models were trained on 8, 16, and 32 A800 GPUs, respectively. This fine-tuning process yields a family of models specialized in website generation, which we name WebGen-LM.
@@ 4.2 Experimental Results
Main Results. Based on the experimental results, we make the following observations: (1) As shown in Tab. 4, WebGen-LM-32B achieves the highest accuracy of 38.2%, surpassing the best proprietary model, DeepSeek-R1, by 10.4%, demonstrating the effectiveness of our training set and the rejection-sampling process. (2) Bolt.diy with DeepSeek-R1 as the engine achieves the highest accuracy among general LLMs at 27.8%, closely followed by Claude-3.5-Sonnet with an accuracy of 26.4%. This indicates that the best-performing models are still far from saturating WebGen-Bench, highlighting that our benchmark remains challenging for current LLMs and agent frameworks. (3) Smaller general open-source models, such as Qwen2.5-Coder-32B and Qwen2.5-72B-Instruct, show a significant performance gap compared to proprietary models. (4) In terms of appearance scores, Bolt.diy with Claude-3.5-Sonnet achieves the best performance of 3.0. The appearance score exhibits a loose correlation with accuracy, as functional webpages typically do not suffer from major rendering issues.
Categorical Results. Apart from the three main instruction categories (shown in Tab. 2), we also classify the test cases into three primary categories based on what they are intended to evaluate: Functional Testing, Data Display Testing, and Design Validation Testing. Detailed definitions and statistics for these categories are provided in Fig. 14 and Tab. 8 in AppendixK. As shown in Tab. 4, among the different categories of test cases, Design Validation Testing achieves the highest accuracy in most cases, while Functional Testing generally yields lower accuracy. Among instruction categories, Content Presentation consistently demonstrates the highest accuracies. This indicates that superficial aspects, such as color themes, are easier to implement than deeper internal functionalities.
@@ 4.3 Ablation Studies
Analysis of the Accuracy of UI Agent Testing Results. To analyze the accuracy of the UI agent testing process, we manually examined three sets of testing results on Bolt.diy. We select the results of Claude-3.5-Sonnet, DeepSeek-R1, and DeepSeek-V3 as the accuracies of these three models are high and are close to each other. The manual testing results serve as the ground truth and require precision; therefore, three human testers independently annotated the results and we assessed the consistency of their annotations. If the annotations of a test case are inconsistent, a fourth human tester is tasked with re-examining the test case and the inconsistent annotations to decide on a final annotation. We present the results of manual testing in Tab. 5. The Alignment Rate is computed with $\text{Alignment Rate}=\frac{N_{\text{Manual}=\text{Agent}}}{N_{\text{total}}}\times 100\%$, where $N_{\text{Manual}=\text{Agent}}$ denotes the number of test cases where the agent-generated result aligns with the manually-annotated result. Further analysis regarding the accuracy of appearance scores is presented in Appendix L.
@@ Appendix D Details of the Decontamination Process
We then inspect whether the remaining training samples contain instructions that are semantic duplicates of the instructions in the testing. For each testing instruction, we retrieve the top-3 training instructions with the highest cosine similarity, and manually inspect them for semantic duplication. We found that the retrieved training samples are all completely different from the testing samples, proving that the final training set is not contaminated. The first three samples in WebGen-Bench and their top matches in the training set are shown in Fig. 5, Fig. 6, and Fig. 7, respectively. The matches are completely different from the test samples.
@@ Appendix I Prompt for Adapting OpenHands Paired with CodeActAgent for WebGen-Bench Evaluation
Appendix K Test Case Categories
@@
Appendix L Analysis of Reliability of Appearance Scores
To strengthen the appearance assessment, we evaluate the bolt.diy results produced by Claude-3.5-Sonnet, DeepSeek-R1, and DeepSeek-V3 with two additional strong multimodal graders—o3 and Claude-3.5-Sonnet. We also manually grade the website screenshots to capture human preference. The results are presented in Tab. 9. Although o3 and Claude-3.5-Sonnet assign slightly lower scores than GPT-4o, the relative ordering of the three results is unchanged (Claude-3.5-Sonnet > DeepSeek-R1 > DeepSeek-V3). Also, for every set of screenshots, the scores descend in the same order: GPT-4o > o3 > Claude-3.5-Sonnet, showing a consistency in the models’ grading pattern. The ensemble results of the three grading models and the human scores both show the same ranking as GPT-4o. These findings indicate that GPT-4o reliably reflects relative appearance quality and aligns with human preference, making it a suitable choice under a limited budget. We will add this discussion to the revised paper.
Table 9: Average appearance scores obtained when the output of each agent–engine LLM is graded by different grading methods.
Agent-engine LLM $\backslash$ Grading method GPT-4o o3 Claude-3.5-Sonnet Average (ensemble) Human
Claude-3.5-Sonnet 3.0 2.7 2.4 2.7 2.8
DeepSeek-R1 2.5 2.3 2.2 2.3 2.3
DeepSeek-V3 2.0 1.9 1.7 1.9 1.9